Please answer this correctly cause there is no point of answering it if you are not sure you just confusing me more
Please answer this correctly cause there is no poi
Answer #1

First, calculate \Delta G_{rxn}:

BrCl\rightarrow \frac{1}{2}Br_{2(g)}+\frac{1}{2}Cl_{2(g)}

\Delta G_{rxn}=\frac{1}{2}\Delta G_{Br_2}+\frac{1}{2}\Delta G_{Cl_2}-\Delta G_{BrCl}

\Delta G_{rxn}=\frac{1}{2}*3.11+\frac{1}{2}*0-(-0.98)=2.535\frac{kJ}{mol}

Now, the relationship between gibbs free energy and the equilibrium constant is:

\Delta G_{rxn}=-RTLn(K_{eq})

So:

2535\frac{J}{mol}=-8.314\frac{J}{mol\cdot K}*298K*Ln(K_{eq})

-1.023=Ln(K_{eq})

K_{eq}=0.3595

Now, we can calculate the concentrations at equilibrium:

\\ BrCl\rightarrow \frac{1}{2}Br_{2(g)}+\frac{1}{2}Cl_{2(g)}\\ 1.3.............0..............0\\ -2x...........x.............x\\ 1.3-2x.......x..............x

K_{eq}=\frac{\frac{x}{12}*\frac{x}{12}}{\frac{1.3-2x}{12}}=0.3595

\frac{x^2}{1.3-2x}=4.314

x^2+8.628x-5.6082=0

x=0.607

So:

n_{BrCl}=1.3-2*0.607=0.086

n_{Br_2}=0.607

n_{Cl_2}=0.607